OpenStudy (anonymous):

Divide eight and two-thirds divided by two and one-third. seven twenty-sixths four and two-thirds three and five-sevenths four and one-half

OpenStudy (jack1):

\[\huge 8 \frac 23~=~8 + \frac 23~ =~ \frac {8 \times 3}{3} + \frac 23~=~\frac {24}{3} + \frac 23\] \[\huge = \frac {26}{3} \]

OpenStudy (jack1):

next 2 and 1 third:

OpenStudy (jack1):

\[\huge 2 \frac 13 = 2 + \frac 13 = \frac {2 \times 3}{3} + \frac 13 = \frac 63 + \frac 13\] \[\huge = \frac 73 \]

OpenStudy (jack1):

next, the operation:

OpenStudy (jack1):

\[ \huge \frac {8 + \frac 23 } { 2 + \frac 13} \] \[ \huge =\frac {\frac {26}{3} } { \frac 73} \] \[ \huge =\frac {26}{3} \times \frac 37 \] ...can u solve from here dude?

OpenStudy (jack1):

@messiahcollins ...? u still there dude?

OpenStudy (anonymous):

yeah

OpenStudy (jack1):

cool, so can u solve it now from the final equation?

OpenStudy (anonymous):

yep

OpenStudy (jack1):

sweet, slaters ;)